# A Novel Testing Approach for Differences Among Brain Connectomes

### v1

@@ 3. Conclusion
Furthermore, $\mathcal{M}$ can be endowed with the structure of a Riemannian manifold by choosing a Riemannian metric. As mentioned before, a popular choice is AIRM (pennec). It can be introduced in a couple of ways. On one hand, it is a Fisher information metric related to the fact that each $C\in\mathcal{M}$ is the set of parameters for a distribution, namely $N\left(0,C\right)$ (lenglet). On the other hand, $\mathcal{M}$ is a homogeneous space of the Lie group $\mathrm{Mat}\left(p\right)^{\times}$ under the action given by $C\mapsto ACA^{T}$ for $A\in\mathrm{Mat}\left(p\right)^{\times}$. The left invariant metric on $\mathrm{Mat}\left(p\right)^{\times}$ descends to a metric on $\mathcal{M}$.
AIRM has a prominent role in part for historic reasons. But it has other nice properties. It is a complete metric: geodesics stay on the manifold for infinite time (the boundary of $\mathcal{M}$, the set of semidefinite matrices, are at infinite distance in AIRM). Also, it has negative curvature. These two facts allow us to use Hadamard’s theorem to conclude that for each $C$, the maps $\exp_{C}:\mathcal{M}\rightarrow T_{C}\mathcal{M}$ and $\log_{C}:T_{C}\mathcal{M}\rightarrow\mathcal{M}$ are global, mutually inverse diffeomorphisms. AIRM introduces a geodesic distance $d\left(\cdot,\cdot\right)$ on $\mathcal{M}$ as well as a norm $\norm{\cdot}_{C}$ on $T_{C}\mathcal{M}$, which satisfy the relation
$\norm{\log_{C}\left(\Lambda\right)}_{C}=d\left(C,\Lambda\right)$
@@
$\mathbb{E}\left[C\right]=\operatorname{argmin}_{\Lambda}\norm{\log_{\Lambda}\left(C\right)}_{\Lambda}^{2}$
@@
Theorem 8 (Jacobi Field Quotient Identity).
Let $M$ be a Riemannian manifold and $\gamma:[0,1]\to M$ be the geodesic $\gamma(s)=\exp_{C}(s\log_{C}(A))$ connecting $C$ to $A$. Consider a variation $\sigma(s,t)$ with:
  • $\sigma(0,t)=C$ for all $t$
  • $\sigma(s,0)=\gamma(s)$
  • $\sigma(1,t)=B(t)$ where $B(0)=B$
Let $J(s)=\frac{\partial\sigma}{\partial t}\big|_{t=0}$ be the associated Jacobi field. Then for small $\varepsilon>0$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}\left[\log_{C}(B)\right]}{\varepsilon}=\nabla_{s}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
@@
Let $\{E_{i}(s)\}_{i=1}^{n}$ be a parallel orthonormal frame along $\gamma$, constructed by parallel transporting an orthonormal basis from $T_{C}M$ along the geodesic. Since the frame is parallel, we have $\nabla_{s}E_{i}=0$ for all $i$ and $s\in[0,1]$.
@@
$\nabla_{s}J=\sum_{i=1}^{n}\frac{dJ^{i}}{ds}E_{i}(s),\quad\nabla_{s}^{2}J=\sum_{i=1}^{n}\frac{d^{2}J^{i}}{ds^{2}}E_{i}(s)$
The Jacobi equation $\nabla_{s}^{2}J+R(J,\dot{\gamma})\dot{\gamma}=0$ becomes:
@@
Note that $\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\big|_{s=0}E_{i}(\varepsilon)$ is the parallel transport of $\nabla_{s}J\big|_{s=0}$ to $T_{\gamma(\varepsilon)}M$.
@@
The covariant derivative $\nabla_{s}J$ along the geodesic has components $\nabla_{s}J]^{i}(s)=\frac{dJ^{i}}{ds}(s)$.
@@
$\nabla_{s}J]^{i}(\varepsilon)=[\nabla_{s}J]^{i}(0)-\varepsilon\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)+O(\varepsilon^{2})$
Therefore, $\nabla_{s}J\big|_{s=\varepsilon}$ expressed in the frame at $\gamma(\varepsilon)$ is:
$\nabla_{s}J\big|_{s=\varepsilon}=\Pi_{C\to\gamma(\varepsilon)}[\nabla_{s}J\big|_{s=0}]-\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
From the quotient expression and the evolution of $\nabla_{s}J$:
$\Pi_{C\to\gamma(\varepsilon)}[\nabla_{s}J\big|_{s=0}]=\nabla_{s}J\big|_{s=\varepsilon}+\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
@@
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\nabla_{s}J\big|_{s=\varepsilon}+\frac{\varepsilon}{2}\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
@@
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\nabla_{s}J\big|_{s=\varepsilon}+O(\varepsilon)$
@@
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)]}{\varepsilon}=\nabla_{s}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
Corollary 9 (Logarithm Difference Formula).
@@
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\nabla_{s}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
where $\tilde{J}(s)$ is the Jacobi field along $\gamma$ with initial conditions:
  • $\tilde{J}(0)=\log_{C}(B)-\log_{C}(A)$
  • $\tilde{J}(1)=0$ (since $\sigma(1,t)=A$ for all $t$ in the variation)
@@
Consider the variation $\sigma(s,t)$ defined by:
This variation satisfies: $\sigma(s,0)=\gamma(s)$ for all $s$, $\sigma(0,t)=\exp_{C}(t[\log_{C}(B)-\log_{C}(A)])$, and $\sigma(1,t)=A$ for all $t$ (since $\gamma(1)=A$ and $\log_{A}(A)=0$).
@@
The first term is $\tilde{J}(\varepsilon)$, which by Theorem 8 equals:
$\tilde{J}(\varepsilon)=\Pi_{C\to\gamma(\varepsilon)}[\tilde{J}(0)]+\varepsilon\nabla_{s}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
@@
$\log_{\gamma(\varepsilon)}(B)-\log_{\gamma(\varepsilon)}(A)=\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]+\varepsilon\nabla_{s}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
@@
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\nabla_{s}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
Lemma 10 (Derivative of Logarithm Map with Moving Base Point).
For a curve $\gamma(s)$ in the manifold of SPD matrices with velocity $\dot{\gamma}(s)$ and a fixed point $B$, the derivative of the logarithm map is:
$\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\nabla_{\dot{\gamma}(0)}\log_{\gamma(0)}(B)$
where $\nabla$ is the Levi-Civita connection of the AIRM metric.
Proof.
This follows from the differential geometry of the logarithm map. The logarithm map $\log:\mathcal{M}\times\mathcal{M}\to T\mathcal{M}$ can be viewed as a map $(P,Q)\mapsto\log_{P}(Q)\in T_{P}\mathcal{M}$. When we vary the base point $P$ along a curve $\gamma(s)$ while keeping $Q=B$ fixed, the change in $\log_{\gamma(s)}(B)$ involves the intrinsic change due to moving the base point (captured by the connection) and the parallel transport of the vector as we move along $\gamma$. The formula $\nabla_{\dot{\gamma}(0)}\log_{\gamma(0)}(B)$ encodes precisely this geometric change. ∎
Proposition 11 (AIRM Jacobi Field Velocity).
@@
$J(s)=\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)$
Then the velocity of the Jacobi field at $s=0$ is given exactly by:
$\dot{J}(0)=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}+\log_{C}(A)$
@@
First, we compute $\frac{d}{ds}\log_{\gamma(s)}(A)$. Since $\gamma(s)$ is the geodesic from $C$ to $A$ with $\gamma(0)=C$ and $\gamma(1)=A$, the remaining distance along the geodesic from $\gamma(s)$ to $A$ is (1s)(1-s) times the initial distance. By the properties of geodesics in Riemannian manifolds:
$\log_{\gamma(s)}(A)=(1-s)\log_{C}(A)$
Therefore:
$\frac{d}{ds}\log_{\gamma(s)}(A)=-\log_{C}(A)$
Next, we compute $\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)$ using Lemma 10. For the AIRM metric at point $C\in\mathcal{M}$, the Levi-Civita connection is:
@@
where $X,Y\in T_{C}\mathcal{M}$ are tangent vectors (symmetric matrices) and $DY[X]$ is the directional derivative.
Applying the lemma with $\dot{\gamma}(0)=\log_{C}(A)$:
$\frac{d}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\nabla_{\log_{C}(A)}\log_{C}(B)$
Since $\log_{C}(A)$ and $\log_{C}(B)$ are constant vector fields when viewed in the tangent space at $C$, the directional derivative term vanishes: $D(\log_{C}(B))[\log_{C}(A)]=0$.
The covariant derivative becomes:
$\nabla_{\log_{C}(A)}\log_{C}(B)=-\frac{1}{2}(C^{-1}\log_{C}(A)\log_{C}(B)+\log_{C}(B)\log_{C}(A)^{T}C^{-1})$
Since $\log_{C}(A)$ and $\log_{C}(B)$ are symmetric matrices, we have $\log_{C}(A)^{T}=\log_{C}(A)$. We can decompose the product of two matrices into symmetric and anti-symmetric parts:
@@
where $\{X,Y\}=XY+YX$ is the anticommutator. After careful computation:
$\nabla_{\log_{C}(A)}\log_{C}(B)=\frac{1}{2}[\log_{C}(A),\log_{C}(B)]+\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}$
Therefore:
@@
Finally, the velocity of the Jacobi field is:
$\dot{J}(0)=\frac{d}{ds}\bigg|_{s=0}[\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)]$
$\left[-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}\right]-(-\log_{C}(A))$
Thus:
$\dot{J}(0)=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}+\log_{C}(A)$
Corollary 12 (Covariant Derivative of Jacobi Field).
Under the same conditions as Proposition 11, the covariant derivative of the Jacobi field at $s=0$ is:
$\nabla_{s}J|_{s=0}=\dot{J}(0)-\frac{1}{2}(C^{-1}\dot{\gamma}(0)J(0)+J(0)\dot{\gamma}(0)^{T}C^{-1})$
where $J(0)=\log_{C}(B)-\log_{C}(A)$ and $\dot{\gamma}(0)=\log_{C}(A)$.
Remark 2.
The formula in Proposition 11 is exact, not an approximation. All terms can be computed using standard matrix operations: matrix logarithms (computable via eigendecomposition), matrix multiplications, and additions. No differential equation solving is required, making the computation both efficient and numerically stable.
The geometric interpretation is clear: the term $\log_{C}(A)$ represents the linear interpolation ensuring $J(1)=0$, the commutator $\log_{C}(A),\log_{C}(B)]$ captures the curvature effect from non-commutativity, and the symmetric term $\{\log_{C}(A),\log_{C}(B)\}_{C}$ arises from the metric structure through the Christoffel symbols.
Tables, Listings and Figures

### v2

@@ A.1 Riemannian Geometry Review
Furthermore, $\mathcal{M}$ can be endowed with the structure of a Riemannian manifold by choosing a Riemannian metric. As mentioned before, a popular choice is AIRM (Pennec, Fillard and Ayache, 2006), which can be introduced in a couple of ways. On one hand, it is a Fisher information metric related to the fact that each $C\in\mathcal{M}$ is the set of parameters for a distribution, namely $N\left(0,C\right)$ (Lenglet et al., 2006b). On the other hand, $\mathcal{M}$ is a homogeneous space of the Lie group $\mathrm{Mat}\left(p\right)^{\times}$ under the action given by $C\mapsto ACA^{T}$ for $A\in\mathrm{Mat}\left(p\right)^{\times}$. The left invariant metric on $\mathrm{Mat}\left(p\right)^{\times}$ descends to a metric on $\mathcal{M}$.
AIRM has a prominent role in and it has several nice properties. It is a complete metric: geodesics stay on the manifold for infinite time (the boundary of $\mathcal{M}$, the set of semidefinite matrices, are at infinite distance in AIRM). Also, it has non-positive curvature. These two facts allow us to use Hadamard’s theorem to conclude that for each $C$, the maps $\exp_{C}:\mathcal{M}\rightarrow T_{C}\mathcal{M}$ and $\log_{C}:T_{C}\mathcal{M}\rightarrow\mathcal{M}$ are global, mutually inverse diffeomorphisms. AIRM introduces a geodesic distance $d\left(\cdot,\cdot\right)$ on $\mathcal{M}$ as well as a norm $\left\lVert\cdot\right\rVert_{C}$ on $T_{C}\mathcal{M}$, which satisfy the relation
$\left\lVert\log_{C}\left(\Lambda\right)\right\rVert_{C}=d\left(C,\Lambda\right)$
@@
$\mathbb{E}\left[C\right]=\operatorname{argmin}_{\Lambda}\left\lVert\log_{\Lambda}\left(C\right)\right\rVert_{\Lambda}^{2}$
@@
A.2 Technical Results
The proof of Theorem 2.4 requires the following estimates.
Theorem A.1 (Jacobi Field Quotient Identity).
Let $M$ be a Riemannian manifold and $\gamma:[0,1]\to M$ be the geodesic $\gamma(s)=\exp_{C}(s\log_{C}(A))$ connecting $C$ to $A$. Consider the variation $\sigma(s,t)=\exp_{\gamma(s)}(t\log(B))$ Let $J(s)=\frac{\partial\sigma(s,t)}{\partial t}\big|_{t=0}$ be the associated Jacobi field. Then for small $\varepsilon>0$:
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}\left[\log_{C}(B)\right]}{\varepsilon}=\frac{D}{ds}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
@@ A.2 Technical Results
Let $\{E_{i}(s)\}_{i=1}^{n}$ be a parallel orthonormal frame along $\gamma$, constructed by parallel transporting an orthonormal basis from $T_{C}M$ along the geodesic. Since the frame is parallel, we have $\frac{D}{ds}E_{i}=0$ for all $i$ and $s\in[0,1]$.
@@
$\frac{D}{ds}J=\sum_{i=1}^{n}\frac{dJ^{i}}{ds}E_{i}(s),\quad\frac{D^{2}}{ds^{2}}J=\sum_{i=1}^{n}\frac{d^{2}J^{i}}{ds^{2}}E_{i}(s)$
The Jacobi equation $\frac{D^{2}}{ds^{2}}J+R(J,\dot{\gamma})\dot{\gamma}=0$ becomes:
@@
Note that $\sum_{i=1}^{n}\frac{dJ^{i}}{ds}\big|_{s=0}E_{i}(\varepsilon)$ is the parallel transport of $\frac{D}{ds}J\big|_{s=0}$ to $T_{\gamma(\varepsilon)}M$.
@@
The covariant derivative $\frac{D}{ds}J$ along the geodesic has components $\frac{D}{ds}J]^{i}(s)=\frac{dJ^{i}}{ds}(s)$.
@@
$\frac{D}{ds}J]^{i}(\varepsilon)=[\frac{D}{ds}J]^{i}(0)-\varepsilon\sum_{j=1}^{n}R_{ij}(0)J^{j}(0)+O(\varepsilon^{2})$
Therefore, $\frac{D}{ds}J\big|_{s=\varepsilon}$ expressed in the frame at $\gamma(\varepsilon)$ is:
$\frac{D}{ds}J\big|_{s=\varepsilon}=\Pi_{C\to\gamma(\varepsilon)}[\frac{D}{ds}J\big|_{s=0}]-\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
From the quotient expression and the evolution of $\frac{D}{ds}J$:
$\Pi_{C\to\gamma(\varepsilon)}[\frac{D}{ds}J\big|_{s=0}]=\frac{D}{ds}J\big|_{s=\varepsilon}+\varepsilon\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
@@
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\frac{D}{ds}J\big|_{s=\varepsilon}+\frac{\varepsilon}{2}\Pi_{C\to\gamma(\varepsilon)}[R(J(0),\dot{\gamma}(0))\dot{\gamma}(0)]+O(\varepsilon^{2})$
@@
$\frac{J(\varepsilon)-\Pi_{C\to\gamma(\varepsilon)}[J(0)]}{\varepsilon}=\frac{D}{ds}J\big|_{s=\varepsilon}+O(\varepsilon)$
@@
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)]}{\varepsilon}=\frac{D}{ds}J\bigg|_{s=\varepsilon}+O(\varepsilon)$
Corollary A.2 (Logarithm Difference Formula).
@@
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\frac{D}{ds}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
where $\tilde{J}(s)$ is the Jacobi field along $\gamma$ associated to the variation $\sigma(s,t)$ defined by:
@@
This variation satisfies: $\sigma(s,0)=\gamma(s)$ for all $s$, $\sigma(0,t)=\exp_{C}(t[\log_{C}(B)-\log_{C}(A)])$, and $\sigma(1,t)=\exp_{A}(t\log_{A}(B))$ for all $t$ (since $\gamma(1)=A$ and $\log_{A}(A)=0$).
@@
The first term is $\tilde{J}(\varepsilon)$, which by Theorem A.1 equals:
$\tilde{J}(\varepsilon)=\Pi_{C\to\gamma(\varepsilon)}[\tilde{J}(0)]+\varepsilon\frac{D}{ds}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
@@
$\log_{\gamma(\varepsilon)}(B)-\log_{\gamma(\varepsilon)}(A)=\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]+\varepsilon\frac{D}{ds}\tilde{J}\big|_{s=\varepsilon}+O(\varepsilon^{2})$
@@
$\frac{\log_{\gamma(\varepsilon)}(B)-\Pi_{C\to\gamma(\varepsilon)}[\log_{C}(B)-\log_{C}(A)]-\log_{\gamma(\varepsilon)}(A)}{\varepsilon}=\frac{D}{ds}\tilde{J}\bigg|_{s=\varepsilon}+O(\varepsilon)$
Proposition A.3 (AIRM Jacobi Field Velocity).
@@
Then
$\frac{D}{ds}J\bigg|_{s=0}=-\frac{1}{2}[\log_{C}(A),\log_{C}(B)]-\frac{1}{2}\{\log_{C}(A),\log_{C}(B)\}_{C}+\log_{C}(A)$
@@
From the definition of $\sigma$,
$J(s)=\log_{\gamma(s)}(B)-\log_{\gamma(s)}(A)$
For the AIRM metric at point $C\in\mathcal{M}$, the Levi-Civita connection is:
@@
where $X,Y\in T_{C}\mathcal{M}$ are tangent vectors (symmetric matrices) and $DY[X]$ is the Euclidean directional derivative.
Then
$\frac{D}{ds}\bigg|_{s=0}\log_{\gamma(s)}(B)=-\frac{1}{2}(C^{-1}\log_{C}(A)\log_{C}(B)+\log_{C}(B)\log_{C}(A)^{T}C^{-1})$
Decomposing the product as:
@@
where $\{X,Y\}=XY+YX$ is the anticommutator. We obtain:
@@
Similarly,
$\frac{D}{ds}\bigg|_{s=0}\log_{\gamma(s)}(A)=-\log_{C}(A)$
And the result follows by gathering the terms.
Remark A.4.
The formula in Proposition A.3 is explicit. All terms can be computed using standard matrix operations: matrix logarithms (computable via eigendecomposition), matrix multiplications, and additions. No differential equation solving is required, making the computation both efficient and numerically stable.
{funding}
This research was supported in part by NIH grant R01 NS112303.